OpenStudy (ksaimouli):

3x^2-4x-2/(x-2)^2(x+2) partial fractions plz help me last one

OpenStudy (ksaimouli):

help

OpenStudy (anonymous):

Factors for 3x^2 - 4x - 2 = ?

OpenStudy (ksaimouli):

no no only bottom one partial fractions

OpenStudy (ksaimouli):

decomposing of fractions

OpenStudy (anonymous):

(x-2)^2(x+2)?

OpenStudy (anonymous):

(x-2)^2 = x^2 - 4x + 4 (x-2)^2 (x+2) = (x^2 - 4x + 4)(x+2) = x^3 - 4x^2 +4x + 2x^2 -8x + 8 = x^3 -2x^2-4x +8

OpenStudy (ksaimouli):

ok i know this i need to find lcd and those stuffs

OpenStudy (ksaimouli):

after fracoring it

OpenStudy (ksaimouli):

if u know plz post it i am laving because i am using 4g internet it costs tooo much

OpenStudy (anonymous):

OK.